Define and plot the percentage error in the kineti
Answer #1

Relativistic:

E_{kr}=mc^{2}-m_{0}c^{2}=m_{0}c^{2}\left ( \frac{1}{\sqrt{1-\left ( \frac{v}{c} \right )^{2}}}-1 \right )

Classic: E_{kc}=\frac{mv^{2}}{2}

Consider Ekr=100%, and the difference between Ek=requested percent.

\frac{E_{kr}-E_{kc}}{E_{kr}}=\frac{c^{2}\left ( \frac{1}{\sqrt{1-\left ( \frac{v}{c} \right )^{2}}} \right )-\frac{v^{2}}{2}}{c^{2}\left ( \frac{1}{\sqrt{1-\left ( \frac{v}{c} \right )^{2}}} \right )}

Not very sure if this is the percent asked in the text (such a parameter may be calculated in many ways).

I choose to work for Ekc/Ekr, which looks like:

\frac{E_{kc}}{E_{kr}}=\frac{x^{2}}{2\left ( \frac{1}{\sqrt{1-x^{2}}}-1 \right )}

x=v/c

The plot of Ekc/Ekr=fc(x) looks like (I might take too many points, but just to be sure the graph looks as it should):

0.4 0.6 0.8

I think the plot of \frac{\Delta E}{E_{kr}}=fc\left ( \frac{v}{c} \right ) looks similar.